# Curriculum Approximate Unlearning for Session-based Recommendation

### v1

@@ Abstract.
Approximate unlearning for session-based recommendation refers to eliminating the influence of specific training samples from the recommender without retraining of (sub-)models. Gradient ascent (GA) is a representative method to conduct approximate unlearning. However, there still exist dual challenges to apply GA for session-based recommendation. On the one hand, naive applying of GA could lead to degradation of recommendation performance. On the other hand, existing studies fail to consider the ordering of unlearning samples when simultaneously processing multiple unlearning requests, leading to sub-optimal recommendation performance and unlearning effect.
To address the above challenges, we introduce CAU, a curriculum approximate unlearning framework tailored to session-based recommendation. CAU handles the unlearning task with a GA term on unlearning samples. Specifically, to address the first challenge, CAU formulates the overall optimization task as a multi-objective optimization problem, where the GA term for unlearning samples is combined with retaining terms for preserving performance. The multi-objective optimization problem is solved through seeking the Pareto-Optimal solution, which achieves effective unlearning with trivial sacrifice on recommendation performance. To tackle the second challenge, CAU adopts a curriculum-based sequence to conduct unlearning on batches of unlearning samples. The key motivation is to perform unlearning from easy samples to harder ones. To this end, CAU first introduces two metrics to measure the unlearning difficulty, including gradient unlearning difficulty and embedding unlearning difficulty. Then, two strategies, hard-sampling and soft-sampling, are proposed to select unlearning samples according to difficulty scores. To verify the effectiveness of the proposed method, we conduct intensive experiments on three benchmark datasets and ablation studies show all the components and strategies of CAU contribute significantly to its superior effectiveness.
@@
submissionid: 452ccs: Information systems Recommender systemsccs: Security and privacy Web application security
@@ 1. Introduction
Session-based recommendation has shown effectiveness in predicting users’ future interests from historical interactions. Despite memorizing sequential patterns, the ability to eliminate the influence of specific training samples, a.k.a. recommendation unlearning (RU), is also crucially important. Several data protection regulations, such as the EU General Data Protection Regulation (GDPR), the California Consumer Privacy Act (CCPA), and the China Personal Information Protection Law (PIPL), have been proposed to ensure individuals’ rights to remove personal data from trained recommendation models (Mantelero, 2013; Baik, 2020). Beyond legal perspective, RU can further serve practical purposes, such as removing outdated user preferences (Liu et al., 2022; Matuszyk et al., 2015) or eliminating the impact of noisy training samples (Shaik et al., 2023).
@@
Exact recommendation unlearning attempts to completely eradicate the impact of the unlearning data by physically removing the unlearning samples and then retraining the model (Nguyen et al., 2022). To improve the unlearning efficiency, existing exact RU methods usually involve data partitioning and train sub-models on each data shard (Chen et al., 2022a, b; Bourtoule et al., 2019). The final results come from the aggregation of sub-models. However, when a large number of unlearning requests occur concurrently, the efficiency of exact RU could also deteriorate substantially, since the unlearning data are likely distributed across multiple shards, necessitating the retraining of several sub-models, or in the worst case, the entire model. Besides, due to prevalent item correlations, such exact RU methods may fail to achieve the exact unlearning effect (Xin et al., 2024). Even if the unlearning interaction has been removed from the data, it can still be inferred from correlated interactions due to the collaborative signals.
@@
Approximate recommendation unlearning offers a more efficient alternative by enabling rapid forgetting and lowering the overall computational cost in contrast. Approximate recommendation unlearning denotes the process of efficiently removing the influence of specific training samples from a recommendation model without retraining the entire model or sub-models from scratch (Li et al., 2024). Gradient ascent (GA) is one of the most representative methods to conduct approximate recommendation unlearning. As deep learning models are typically trained using gradient descent-based optimization, an intuitive approach for unlearning is to get the opposite gradient for unlearning samples, and then add the gradient back to increase the loss function. GA mitigates the influence of unlearning samples from a parametric perspective, achieving the approximate unlearning effect. Yao et al. (2024) proposed to utilize GA-based methods to conduct unlearning for large language models.
@@
Existing challenges. Despite its success, directly applying GA to conduct unlearning for session-based recommendation still poses non-trivial challenges.
i). Simply applying GA for session-based recommendation could result in significant degradation of recommendation performance. Since the objective function of GA is an inverted version of gradient descent, the objective function would diverge without tailored strategies, leading to failure of the entire training. Besides, due to the sequential correlations in session-based recommendation, conducting GA in one session could negatively impact the accuracy of other sessions. Thus, achieving a proper trade-off between unlearning and retaining through GA presents a key research challenge.
@@
ii). Existing research fails to consider the sequence of unlearning samples when handling a batch of unlearning requests simultaneously. Curriculum learning has shown effectiveness in improving the learning process through learning from easy samples to hard ones. However, the effect of unlearning orders remains underexplored for session-based recommendation when handling a batch of unlearning requests. Intuitively, due to sequential correlations between interactions, the ordering of unlearning instances could significantly impact the unlearning outcome.
@@
The proposed method. In this paper, we propose curriculum approximate unlearning (CAU), an approximate unlearning framework tailored to session-based recommendation. Concretely, the unlearning task is formulated as a GA term on unlearning samples. To address the first challenge, CAU defines the overall optimization task as a multi-objective optimization problem, where the GA term for unlearning samples is combined with retaining terms which aim to preserve the recommendation performance on remaining samples. Through seeking the Pareto-Optimal solution, CAU achieves effective unlearning with trivial sacrifice.
@@
To address the second challenge, CAU uses a curriculum-based sequence to conduct unlearning on batches of unlearning samples. More precisely, we first propose two metrics to quantify the unlearning difficulty, including the gradient unlearning difficulty and embedding unlearning difficulty. The gradient unlearning difficulty reflects the consistency between the GA term gradient and the retaining term gradient, while the embedding unlearning difficulty captures semantic proximity between session states and unlearning items. Then, CAU introduces curriculum-based strategies to conduct unlearning according to the difficulty scores. Specifically, CAU introduces two unlearning strategies: (1) hard-sampling: a deterministic curriculum-based approach that strictly follows the unlearning difficulty ranking to select samples in every unlearning epoch, and (2) soft-sampling: a stochastic sampling strategy that selects easier samples with higher probability in early steps and gradually shifts towards harder ones in later steps.
@@
To verify the effectiveness of the proposed method, CAU is implemented on three representative session-based recommendation models, including GRU4Rec, SASRec, and BERT4Rec. Extensive experiments on three benchmark datasets are conducted, and experimental results show the effectiveness of the proposed method.
Contributions. To summarize, the main contributions lie in:
@@
  • To the best of our knowledge, CAU is the first attempt that conducts approximate unlearning for session-based recommendation through curriculum-based multi-task optimization. Pareto-Optimization is utilized to achieve the trade-off between unlearning and retaining recommendation performance.
@@
  • We propose gradient unlearning difficulty and embedding unlearning difficulty to quantify unlearning difficulty in the context of session-based recommendation.
  • We propose two curriculum learning strategies to guide the unlearning process, starting with easier samples and progressively moving toward more difficult ones.
@@
  • We conduct extensive experiments on three state-of-the-art session-based recommendation models and three benchmark datasets to show that CAU achieves efficient and effective unlearning while retaining high recommendation performance.
@@ 2.1. Session-based recommendation
Session-based recommendation aims to predict a user’s next action based on the current session. Early methods rely on session-based nearest-neighbor techniques, which recommend items frequently co-occurring with the current session  (Hariri et al., 2012, 2015; Bell and Koren, 2007). These methods are simple and interpretable but struggle to capture complex sequential patterns. With the rise of deep learning, RNN-based models  (Donkers et al., 2017; Hidasi and Karatzoglou, 2018; Mu, 2018) were proposed to model session sequences, achieving better performance by capturing temporal dependencies. Later works explored self-attention and Transformer-based architectures to model long-range dependencies and complex item interactions within sessions  (Kang and McAuley, 2018; Sun et al., 2019; Xu et al., 2019; Fang, 2021). Recent research further incorporates graph neural networks (GNNs) to represent sessions as graphs  (Wu et al., 2019; Wang et al., 2019; Qiu et al., 2019), capturing intricate item transition patterns and higher-order connectivity. Beyond model architecture, there are efforts on improving training and personalization strategies, including contrastive learning  (Liu et al., 2021; Xie et al., 2022), and meta-learning  (Song et al., 2021; Pan et al., 2022) for session-based recommendation. These approaches aim to better exploit session context, mitigate data sparsity, and improve generalization to unseen sessions.
@@ 2.3. Recommendation unlearning
Unlearning in the recommendation scenario tends to attract more research attention. Recommendation unlearning can not only help to protect user privacy but also improve recommendation models through eliminating the effect of noisy data and misleading information  (Shaik et al., 2023).  Liu et al. (2022) and  Xu et al. (2023) proposed to use fine-tuning and the alternative least square algorithm for unlearning acceleration.  Chen et al. (2022a) and  Li et al. (2022) extended the ideas of the SISA algorithm for collaborative filtering. And  Hu et al. (2025) applied the similar idea to LLM-based recommendation unlearning.  Xin et al. (2024) proposed an exact unlearning framework for session-based recommendation. Additionally, there are also other techniques like matrix correction  (Liu et al., 2023), influence function  (Zhang et al., 2024) and label-flipping  (Alshehri and Zhang, 2023). However, none of the existing approximate unlearning methods is tailored for session-based recommendation. Previous works also do not take into account the difficulty of the unlearning samples, which may lead to suboptimal performance. In this work, we propose two methods to estimate the difficulty of unlearning samples and introduce two corresponding sampling strategies to implement a curriculum learning scheme, thereby further improving the effectiveness of the unlearning process.
@@ 3.1. Session-based recommendation
Let $V$ be the set of items, $S=\{v_{1},...,v_{t},...,v_{n}\}$ be the user-item interaction sequence in a session, where $v_{t}\in V$ is the item interacted at time step $t$ in the session. The dataset $D$ of user-item interaction sessions is defined as $D=\{S_{1},...,S_{|D|}\}$. Given a user-item interaction sequence at time step $t$, the learning objective of session-based recommendation is to predict the next item $v_{t+1}$ that the user is likely to interact with, which can be formalized as:
(1) $\mathbf{p}_{t+1}=\mathcal{M}_{\theta}(v|S,D)$
where $\mathbf{p}_{t+1}$ denotes the predicted probability distribution over candidate items at time step $t+1$, $\mathcal{M}_{\theta}$ represents the recommendation model parameterized by $\theta$, e.g., GRU4Rec (Hidasi et al., 2015) and SASRec (Kang and McAuley, 2018). Typically, the recommendation model $\mathcal{M}_{\theta}$ can be trained by optimizing a loss function $\mathcal{L}_{rec}$ based on the predicted probability distribution $\mathbf{p}_{t+1}$, such as the cross-entropy loss or the pair-wise ranking loss. The training process can be formalized as:
(2) $\theta^{rec}=\arg\min_{\theta}\mathcal{L}_{rec}(\theta^{inital};D)$
@@
Figure 1. Overview of the proposed CAU framework. CAU defines the unlearning objective $\mathcal{L}_{unlearn}$ and remaining terms $\mathcal{L}_{normal}$ and $\mathcal{L}_{KL}$. Then CAU introduces two metrics to measure the unlearning difficulty, including gradient difficulty and embedding difficulty. Two sampling methods are proposed to sample batches from easy to hard, according to the unlearning difficulty metrics. Finally, a Pareto-Optimality solver is used to perform multi-objective learning to obtain the unlearned model.
@@ 3.2. Unlearning
Exact recommendation unlearning targets on completely eradicating the impact of the data to be forgotten as if they never occurred in the training process. Let $D_{f}\subset D$ denotes the data to be unlearned, the naive approach to achieve exact unlearning is to retrain the model from scratch on the dataset without the unlearning data, i.e., $D-D_{f}$. This process can be defined as:
(3) $\theta^{exa}=\arg\min_{\theta}\mathcal{L}_{rec}(\theta^{inital};D-D_{f})$
@@
In contrast to exact unlearning, approximate recommendation unlearning offers a more computationally efficient alternative that aims to emulate the effect of exact unlearning, while preserving model utility and avoiding retraining. Instead of retraining the model from scratch, approximate unlearning optimizes the model parameters directly in response to unlearning requests. Common methods for approximate unlearning include gradient ascent on unlearning samples to reverse their influence (Graves et al., 2021; Sekhari et al., 2021) or filtering the model’s outputs (Baumhauer et al., 2022). Approximate recommendation unlearning can be defined as:
(4) $\theta^{app}=\arg\min_{\theta}\mathcal{L}_{app}(\theta^{rec};D_{f})$
@@ 4. Method
The framework is illustrated in Figure 1. Specifically, to maintain recommendation performance after unlearning, CAU formulates the task as a multi-objective learning problem. CAU incorporates a learning objective that focuses on preserving recommendation performance, alongside the unlearning objective for the unlearning data $D_{f}$. We then delve into the model’s gradient and embedding space and introduce two metrics to measure unlearning difficulty, respectively. Based on these difficulty metrics, CAU proposes two sampling methods—hard-sampling and soft-sampling—to construct unlearning batches, which facilitates a curriculum strategy to perform unlearning from easy to hard. Finally, a Pareto-Optimality solver is utilized to achieve multi-objective optimization to obtain the unlearned model parameters.

### v2

@@ Abstract.
Approximate unlearning aims to eliminate the influence of specific training samples from a recommender system without the need for model retraining. Gradient ascent (GA) is a prominent method for approximate unlearning but applying it to session-based recommendation presents two primary challenges. First, a naive application of GA often results in a significant degradation of recommendation performance. Second, existing studies fail to account for the sequential ordering of unlearning samples when processing multiple simultaneously unlearning requests, leading to sub-optimal efficacy and recommendation performance. To address these challenges, we introduce CAU, a curriculum approximate unlearning framework tailored to session-based recommendation. CAU executes unlearning by applying a GA term to the target unlearning samples. To mitigate performance degradation (the first challenge), CAU treats unlearning as a multi-objective optimization problem. By integrating the GA term with retaining terms for preserving performance, the framework identifies a Pareto-Optimal solution. This approach ensures robust unlearning with minimal impact on recommendation performance. To resolve the second challenge regarding sample ordering, CAU adopts a curriculum-based sequence to process unlearning batches. Motivated by the principle of progressing from “easy” to “hard” samples, we introduce two metrics to quantify unlearning difficulty: gradient difficulty and embedding difficulty. We further propose two selection strategies, hard-sampling and soft-sampling, to prioritize unlearning samples on these difficulty scores. Extensive experiments on three benchmark datasets, alongside comprehensive ablation studies, demonstrate that each component of CAU contributes significantly to its superior performance and unlearning efficacy.
@@
$c$s: Information systems Recommender systemsccs: Security and privacy Web application security
@@ 1. Introduction
Session-based recommendation has shown remarkable effectiveness in predicting users’ future interests based on their historical interactions. While these systems excel at memorizing sequential patterns, the ability to eliminate the influence of specific training samples, a.k.a. recommendation unlearning (RU), has emerged as equally crucial. Several data protection regulations, such as the EU General Data Protection Regulation (GDPR), the California Consumer Privacy Act (CCPA), and the China Personal Information Protection Law (PIPL), mandate individuals’ rights to have their personal data removed from trained recommendation models (Mantelero, 2013; Baik, 2020; Schelter et al., 2023). Beyond this legal perspective, RU serves additional practical purposes, such as removing outdated user preferences (Liu et al., 2022; Matuszyk et al., 2015) or eliminating the impact of noisy training samples (Shaik et al., 2023; Schelter et al., 2024).
@@
Exact recommendation unlearning eradicates unlearning data impact by removing samples and retraining the model (Nguyen et al., 2022). To improve efficiency, exact RU methods typically partition data and train sub-models on each shard (Chen et al., 2022a, b; Bourtoule et al., 2019), then aggregate results. However, when numerous unlearning requests occur concurrently, efficiency deteriorates substantially as unlearning data spans multiple shards, requiring retraining of several sub-models or the entire model. Moreover, due to item correlations, exact RU methods may fail to achieve true unlearning (Xin et al., 2024), as removed interactions can still be inferred from correlated interactions through collaborative signals.
@@
Approximate recommendation unlearning offers a more efficient alternative by enabling rapid forgetting and lowering overall computational costs. It efficiently removes the influence of specific training samples from a recommendation model without retraining the entire model or sub-models from scratch (Li et al., 2024). Gradient ascent (GA) is a representative method that applies opposite gradients to unlearning samples, increasing the loss function to mitigate their influence from a parametric perspective. Yao et al. (2024) proposed to use GA-based methods for unlearning in large language models.
@@
Challenges. Despite its success, directly applying GA to conduct unlearning for session-based recommendation still poses non-trivial challenges:
  1. (1)
    Naive application of GA for session-based recommendation can severely degrade performance. Because GA’s objective function is the inverse of gradient descent, it may diverge without specific strategies, causing training failure. Furthermore, due to sequential correlations, performing GA on one session can negatively impact the accuracy of others. Consequently, balancing unlearning efficacy with maintaining performance via GA remains a significant challenge.
  2. (2)
@@
    Existing research overlooks the sequence of unlearning samples when processing simultaneous requests. While curriculum learning effectively improves training by progressing from easy to hard samples, the impact of unlearning orders remains underexplored for session-based recommendation. Intuitively, sequential correlations between interactions suggest that the ordering of unlearning instances significantly influences the final outcome.
@@
Our proposed method. We propose curriculum approximate unlearning (CAU), a framework for session-based recommendation that uses gradient ascent (GA) to remove data influence. To address the first challenge and maintain recommendation quality, CAU formulates unlearning as a multi-objective optimization problem, balancing the GA term with performance-retention objectives. By identifying a Pareto-optimal solution, CAU achieves effective unlearning with minimal performance loss.
@@
To address the second challenge, CAU employs a curriculum-based sequence to process unlearning batches. We quantify difficulty using two metrics: gradient unlearning difficulty, which measures the consistency between GA and retaining gradients, and embedding unlearning difficulty, which captures the semantic proximity between session states and target items. Based on these scores, CAU implements two strategies: hard-sampling, a deterministic approach strictly following difficulty rankings, and soft-sampling, a stochastic approach that prioritizes easier samples early on before gradually shifting toward harder ones.
@@
To verify CAU’s effectiveness, we implement it on three representative session-based recommendation models, viz. GRU4Rec, SASRec, and BERT4Rec. We conduct extensive experiments on three benchmark datasets. Our experimental results show CAU’s effectiveness.
Contributions. To summarize, our main contributions in this paper are:
@@
  • We propose CAU, the first curriculum-based multi-task optimization framework for approximate unlearning in session-based recommendation, using Pareto-optimization to balance unlearning efficacy with performance retention.
@@
  • We introduce gradient and embedding unlearning difficulty metrics to quantify the complexity of removing specific samples in session-based contexts.
  • We design two curriculum learning strategies (hard-sampling and soft-sampling) that optimize the unlearning process by progressing from easy to difficult samples.
@@
  • Extensive experiments across three benchmark datasets and three state-of-the-art models demonstrate that CAU achieves superior unlearning efficiency and effectiveness while maintaining high recommendation accuracy.
@@ 2.1. Session-based recommendation
Session-based recommendation aims to predict a user’s next action based on the current session. Early methods rely on session-based nearest-neighbor techniques, which recommend items frequently co-occurring with the current session (Hariri et al., 2012, 2015; Bell and Koren, 2007). These methods are simple and interpretable but struggle to capture complex sequential patterns.
With the rise of deep learning, RNN-based models (Donkers et al., 2017; Hidasi and Karatzoglou, 2018; Mu, 2018) were proposed to model session sequences, achieving better performance by capturing temporal dependencies. Later works explored self-attention and Transformer-based architectures to model long-range dependencies and complex item interactions within sessions (Kang and McAuley, 2018; Sun et al., 2019; Xu et al., 2019; Fang, 2021). Recent research further incorporates graph neural networks (GNNs) to represent sessions as graphs (Wu et al., 2019; Wang et al., 2019; Qiu et al., 2019), capturing intricate item transition patterns and higher-order connectivity. Beyond model architecture, there are efforts on improving training and personalization strategies, including contrastive learning (Liu et al., 2021; Xie et al., 2022), and meta-learning (Song et al., 2021; Pan et al., 2022). These approaches aim to better exploit session context, mitigate data sparsity, and improve generalization to unseen sessions.
@@ 2.2. Machine unlearning
While several contemporary studies have adapted multi-task optimization (MTO) to LLM unlearning (Jin et al., 2025; Pan et al., 2025), it is crucial to note that recommendation tasks are fundamentally distinct from general LLM-based tasks. Specifically, recommendation systems primarily rely on collaborative signals and user-item interaction graphs, which exhibit higher data sparsity compared to the dense semantic spaces of LLMs. Consequently, directly migrating MTO frameworks from LLMs to recommendation unlearning without considering these domain-specific characteristics may lead to suboptimal performance and the loss of critical collaborative information.
@@ 2.3. Recommendation unlearning
Unlearning in the recommendation scenario tends to attract more research attention. Recommendation unlearning can not only help to protect user privacy but also improve recommendation models through eliminating the effect of noisy data and misleading information (Shaik et al., 2023; Lubitzsch et al., 2025). Liu et al. (2022) and Xu et al. (2023) proposed to use fine-tuning and the alternative least square algorithm for unlearning acceleration. Chen et al. (2022a) and Li et al. (2022) extended the ideas of the SISA algorithm for collaborative filtering. And Hu et al. (2025) applied a similar idea to LLM-based recommendation unlearning. Xin et al. (2024) proposed an exact unlearning framework for session-based recommendation. Additionally, there are other techniques like matrix correction (Liu et al., 2023), influence functions (Zhang et al., 2024), and label-flipping (Alshehri and Zhang, 2023). However, none of the existing approximate unlearning methods is tailored for session-based recommendation. Recently, Dou et al. (2025) introduced interaction-level unlearning difficulty metrics for collaborative filtering to guide extra deletion but does not leverage difficulty to optimize the unlearning process. In this work, we propose two methods to estimate the difficulty of unlearning samples and introduce two corresponding sampling strategies to implement a curriculum learning scheme, thereby further improving the effectiveness of the unlearning process.
@@ 3.1. Session-based recommendation
Let $V$ be the set of items, $S=\{v_{1},\ldots,v_{t},\ldots,v_{n}\}$ be the user-item interaction sequence in a session, where $v_{t}\in V$ is the item interacted at time step $t$ in the session. The dataset $D$ of user-item interaction sessions is defined as $D=\{S_{1},\ldots,S_{|D|}\}$. Given a user-item interaction sequence at time step $t$, the learning objective of session-based recommendation is to predict the next item $v_{t+1}$ that the user is likely to interact with, which can be formalized as:
(1) $\mathbf{p}_{t+1}=\mathcal{M}_{\theta}(v|S,D),$
where $\mathbf{p}_{t+1}$ denotes the predicted probability distribution over candidate items at time step $t+1$, $\mathcal{M}_{\theta}$ represents the recommendation model parameterized by $\theta$, e.g., GRU4Rec (Hidasi et al., 2015) or SASRec (Kang and McAuley, 2018). Typically, the recommendation model $\mathcal{M}_{\theta}$ can be trained by optimizing a loss function $\mathcal{L}_{rec}$ based on the predicted probability distribution $\mathbf{p}_{t+1}$, such as the cross-entropy loss or the pair-wise ranking loss. The training process can be formalized as:
(2) $\theta^{rec}=\arg\min_{\theta}\mathcal{L}_{rec}(\theta^{inital};D),$
@@
Figure 1. Overview of the proposed CAU framework. CAU defines an unlearning objective $\mathcal{L}_{unlearn}$ and remaining terms $\mathcal{L}_{normal}$ and $\mathcal{L}_{KL}$. Then CAU introduces two metrics to measure the unlearning difficulty, including gradient difficulty and embedding difficulty. Two sampling methods are proposed to sample batches from easy to hard, according to the unlearning difficulty metrics. Finally, a Pareto-optimality solver is used to perform multi-objective learning to obtain the unlearned model.
@@ 3.2. Unlearning
Exact recommendation unlearning is aimed at completely eradicating the impact of the data to be forgotten as if they never occurred in the training process. Let $D_{f}\subset D$ denotes the data to be unlearned, the naive approach to achieve exact unlearning is to retrain the model from scratch on the dataset without the unlearning data, i.e., $D-D_{f}$. This process can be defined as:
(3) $\theta^{exa}=\arg\min_{\theta}\mathcal{L}_{rec}(\theta^{inital};D-D_{f}),$
@@
In contrast to exact unlearning, approximate recommendation unlearning offers a computationally more efficient alternative that aims to emulate the effect of exact unlearning, while preserving model utility and avoiding retraining. Instead of retraining the model from scratch, approximate unlearning optimizes the model parameters directly in response to unlearning requests. Common methods for approximate unlearning include gradient ascent on unlearning samples to reverse their influence (Graves et al., 2021; Sekhari et al., 2021) or filtering the model’s outputs (Baumhauer et al., 2022). Approximate recommendation unlearning can be defined as:
(4) $\theta^{app}=\arg\min_{\theta}\mathcal{L}_{app}(\theta^{rec};D_{f}),$
@@ 4. Method
The framework is illustrated in Figure 1. To maintain recommendation performance after unlearning, CAU formulates the task as a multi-objective learning problem. CAU incorporates a learning objective that focuses on preserving recommendation performance, alongside the unlearning objective for the unlearning data $D_{f}$. We then consider the model’s gradient and embedding space and introduce two metrics to measure unlearning difficulty, respectively. Based on these difficulty metrics, CAU proposes two sampling methods, hard-sampling and soft-sampling, to construct unlearning batches, which facilitates a curriculum strategy to perform unlearning from easy to hard. Finally, a Pareto-optimality solver is used to achieve multi-objective optimization to obtain the unlearned model parameters.